A billiard ball moving at a speed of 7.95 m/s strikes an identical stationary ball a glancing blow. After the collision, one ball is found to be moving at a speed of 1.80 m/s in a direction making a 59.5 ° with the original line of motion. What is the speed of the other ball? At what angle is it moving? Give your answer in degrees. (Hint: use conservation of linear momentum)
Answer #1

let be m the mass of the ball:

the intial linear momentum is:

p_o = 7.95 m this initial momentum is along x axis, there is no y component of the linear momentum:

let be \theta the angle between the diretion of the velocity of the second mass and the intial direction.

In the horizontal axis:

7.95 m = m(1.8\cos(59.5))+mv_2\cos\theta

v_2\cos\theta = 7.036......1

in the vertical axis:

p_y_1=p_y_2

m(1.8\sin(59.5)) = mv_2\sin\theta

v_2\sin\theta = 1.55.....2

from 1 and 2:

\theta =12.4

and:

v_2 = 7.218 m/s